OpenStudy (anonymous):

Find the x and y values at the minimum of 2x^2+xy+2y^2+3y+2 And find the x and y values at the maximum of -3x^2+2xy-4y^2+2y+7

OpenStudy (jonnyvonny):

Are you in calculus?

OpenStudy (anonymous):

not really

OpenStudy (jonnyvonny):

How "not really"?... I'd imagine either yes or no. Lol.

OpenStudy (jonnyvonny):

Do you know how to take the derivative?

OpenStudy (anonymous):

its a calculus class geared towards the social sciences we haven't covered this material

OpenStudy (jonnyvonny):

Mmk, so you don't know how to take the derivative? Or do you?

OpenStudy (anonymous):

i dont is it ok if you can just go through the whole thing, and i'll teach myself from there

OpenStudy (jonnyvonny):

Well, in which case, you'll be learning 3 different things in this problem; the taking the derivative, the implicit differentiation, and the product rule. Taking a derivative of anything is easy if its by itself, and an x/constant. What you do, is take the power's value, multiply it by the coefficient, and subtract the power by 1. Examples of derivatives: \[5x^1\rightarrow 5x ^{1-1}\rightarrow5x^0=5\] \[3x^2\rightarrow 6x ^{2-1}=6x^1=6x\]